### v1

@@ 1 Introduction
The holographic principle Maldacena:1997re asserts that a gravitational theory defined in $d$ dimensions can be reformulated through an equivalent boundary theory living in $(d{-}1)$ dimensions. This paradigm inspired investigations into two–dimensional $\mathrm{AdS}_{2}$ holography in the late 1990s, primarily driven by the analysis of asymptotic symmetries in Jackiw–Teitelboim (JT) gravity Teitelboim:1983ux ; Jackiw:1984je and dilaton models arising from dimensional reduction Strominger:1998yg ; Cadoni:1998sg ; Hotta:1998iq ; Navarro-Salas:1999zer . These studies have recently experienced a resurgence, motivated by the conjectured duality between JT gravity and the Sachdev–Ye–Kitaev (SYK) model Maldacena:2016hyu ; Jensen:2016pah ; Sachdev:1992fk ; Kitaev , which exemplifies a solvable system with quantum chaotic behavior. Within this correspondence, the Schwarzian boundary action emerges as a low-energy effective description Mertens:2018fds .
Conformal field theory has emerged as a foundational element in contemporary theoretical physics, with far–reaching applications from string theory to the study of complex many–body systems. Its significance lies in the stringent symmetry principles it enforces, which serve to tightly constrain the landscape of admissible physical models. In two dimensions, conformal invariance proves exceptionally potent, manifesting in the form of the infinite–dimensional Virasoro algebra Virasoro:1969zu . Over time, extensive research has been directed toward extending this symmetry to more generalized algebraic settings, such as the Kac–Moody algebras Kac_1968 ; Moody:1968zz and $W$–algebras Zamolodchikov:1985wn . Simultaneously, substantial effort has been invested in deepening the theoretical understanding of the Virasoro structure itself DiFrancesco:1997nk ; Blumenhagen:2009zz ; Prochazka:2014gqa ; Ozer:2015dha . These algebraic symmetries also underpin the formulation of two–dimensional dilaton gravity, wherein the appearance of the Virasoro algebra at the boundary offers a natural bridge between conformal techniques and gravitational dynamics.
@@
The intricate internal structure of JT gravity is fundamentally grounded in its boundary symmetry content and the formal mathematical scaffolding that supports it. The theory maintains a profound connection to conformal symmetry through its interplay with the Virasoro algebra and Schwarzian derivatives Iliesiu ; Mertens ; OvisenkoSchwarzian , and it has been systematically broadened by the implementation of relaxed boundary conditions and higher–spin extensions Gross:2016kjj ; Witten:2016iux ; Yoon . The boundary behavior can also be effectively captured using coadjoint orbit methods, which serve as powerful tools for probing symmetry–breaking phenomena and their physical ramifications Davison:2016ngz ; NarayanYoon ; Fu:2016vas ; Grumiller:2017qao ; Alkalaev:2013fsa ; Grumiller:2013swa ; Grumiller ; Kine ; Fukuyama . While the Schwarzian formalism remains a crucial link to the SYK dual, the scope of asymptotic symmetry analysis extends beyond it; in particular, the BF formulation of JT gravity offers a structurally independent yet equally robust avenue for such inquiries Grumiller:2002nm ; Banks:1990mk . Within this formalism, the dilaton field functions as a Lagrange multiplier that enforces curvature constraints, and the associated asymptotic symmetry algebra can be extended, for instance, to a Virasoro–Kac–Moody structure, via the inclusion of internal symmetry sectors Ikeda:1993fh ; Jackiw:1982hg . This flexibility underscores the utility of JT gravity as a versatile platform for studying gravitational dynamics in lower dimensions, irrespective of adherence to the Schwarzian perspective.
@@ 2.1 Conformal structure of the quantum ${\cal N}=1$ superconformal algebra
2.2 Extended ${\cal N}=1$ superconformal algebra with conformal spin–$(\Delta,1-\Delta)$
In this section, we construct an extended version of the quantum ${\cal N}=1$ superconformal algebra by supplementing the conformal fields of spin $\Delta=2$ and $\Delta=3/2$, introduced in the previous subsection, with two additional components. These new fields, labeled $X$ and $Y$, are interpreted as dilaton partners and carry conformal weights $11-\Delta=-$ 1 and $-1/2$, respectively. Their standard mode expansions are given by $X(z)=\sum_{n}X_{n}z^{-n+1}$ and $F(z)=\sum_{n}F_{n}z^{-n+1/2}$. Using Thielemans’ Mathematica package Thielemans:1991uw , one obtains the non–vanishing operator product expansions that define the extended ${\cal N}=1$ superconformal algebra in this formulation.
@@ 3.2 Connection to $sl(2,\mathbb{R})$ BF theory
Analogous to the computational benefits afforded by Chern–Simons theory in three–dimensional gravity Achucarro:1986uwr ; Witten:1988hc , alternative formulations in two dimensions offer substantial simplifications. In this work, we briefly summarize the key aspects of such approaches, with particular emphasis on the JT model. Specifically, in two dimensions, dilaton gravity with a negative cosmological constant admits an equivalent description in terms of a gauge theory defined over a spacetime manifold ${\cal M}$:
@@
Here, $\mathcal{A}$ denotes a one–form connection with corresponding field strength ${\cal F}=\mathrm{d}{\cal A}+{\cal A}\wedge{\cal A}$, and $k$ is a coupling constant. The dilaton field ${\cal X}$ is treated as a scalar taking values in the gauge algebra. The spacetime manifold ${\cal M}$ is assumed to have the topology of a cylinder, ${\cal M}\cong S^{1}\times\mathbb{R}_{+}$, where the radial coordinate $\rho$ ranges over $0\leq\rho<\infty$. The Euclidean time coordinate $y$ is periodically identified as $y\sim y+\beta$, with $\beta$ denoting the inverse temperature. A boundary term $S_{\text{bdy}}$ is included to ensure a well–defined variational principle, enforcing appropriate conditions at the asymptotic boundary $\partial\mathcal{M}$ located at infinite radius. The notational framework adopted in this work, including the components of the gauge fields, the dilaton variables, and the boundary charges, is based on standard conventions widely used in the literature. For the sake of clarity and consistency, the principal symbols and their associated physical roles are summarized in Appendix A.
@@
$\operatorname{d}\!s^{2}=\operatorname{d}\!\rho^{2}+2\,\mathcal{L}^{0}\operatorname{d}\!\rho\operatorname{d}\!\varphi+\Big{(}\left(e^{\rho}\mathcal{L}^{+}-e^{-\rho}\mathcal{L}^{-}\right)^{2}+\left(\mathcal{L}^{0}\right)^{2}\Big{)}\,\operatorname{d}\!\varphi^{2}\leavevmode\nobreak\ ,$ (18)
@@
$X=e^{\rho}\,\mathcal{X}^{+}+e^{-\rho}\,\mathcal{X}^{-}\leavevmode\nobreak\ .$ (19)
@@ 3.3 Affine boundary conditions
for any generic phase space functional $\digamma$. The asymptotic symmetry algebra is thus generated by the boundary charges $\mathcal{L}^{i}$ and $\mathcal{X}^{i}$. Ultimately, the corresponding operator product algebra takes the form:
@@
$\mathfrak{\mathcal{\mathfrak{J}}}^{A}(\tau_{1})\mathfrak{\mathcal{\mathfrak{J}}}^{B}(\tau_{2})$ $\sim$ $\frac{\frac{k}{2}\eta^{AB}}{\tau_{12}^{2}}+\frac{\mathfrak{\mathcal{\mathfrak{f}}}^{AB}_{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ C}\mathfrak{\mathcal{\mathfrak{J}}}^{C}}{\tau_{12}}.$ (33)
Here, $\eta^{AB}$ denotes the trace matrix, and $\mathfrak{f}^{AB}_{\leavevmode\nobreak\ \leavevmode\nobreak\ C}$ are the structure constants of the underlying algebra, with indices running over $(A,B=0,\pm 1)$. In particular, one has $\mathfrak{f}^{ij}_{\leavevmode\nobreak\ \leavevmode\nobreak\ i+j}=(i-j)$. Accordingly, under the most relaxed set of boundary conditions for $sl(2,\mathbb{R})$ dilaton gravity, the resulting asymptotic symmetry algebra is identified with a single copy of the affine $\mathfrak{sl}(2,\mathbb{R})_{k}$ algebra.
@@ 3.4 Conformal boundary conditions
Building upon the canonical analysis of the $\mathcal{N}=1$ $\mathfrak{osp}(1|2)$ case, we now formulate the two–dimensional higher-spin dilaton supergravity theory with negative cosmological constant in the BF framework. This theory can be equivalently described by the following action on a spacetime manifold ${\cal M}$:
@@ 4.1 $\mathcal{N}=1$ affine boundary conditions
$\delta_{\lambda}\mathcal{G}^{\mp\frac{1}{2}}$ =±$=\pm k\partial_{t}\eta^{\mp\frac{1}{2}}+2\eta^{\pm\frac{1}{2}}\mathcal{L}^{\mp 1}+2\eta^{\mp\frac{1}{2}}\mathcal{L}^{0}-\mathcal{G}^{\pm\frac{1}{2}}\epsilon^{\mp 1}\mp\frac{1}{2}\mathcal{G}^{\mp\frac{1}{2}}\epsilon^{0},$ (53)
@@
$\delta_{\lambda}\mathcal{Y}^{\mp\frac{1}{2}}$ =±$=\pm\frac{1}{2}\eta^{\mp\frac{1}{2}}\mathcal{X}^{0}\mp\eta^{\pm\frac{1}{2}}\mathcal{X}^{\mp 1}\pm\mathcal{Y}^{\pm\frac{1}{2}}\epsilon^{\mp 1}\mp\frac{1}{2}\mathcal{Y}^{\mp\frac{1}{2}}\epsilon^{0}.$ (56)
@@
for an arbitrary phase space functional $\digamma$. In this setting, the charges $\mathcal{L}^{i}$ and $\mathcal{G}^{i}$, along with their corresponding dilaton partners $\mathcal{X}^{i}$ and $\mathcal{Y}^{i}$, serve as the generators of the asymptotic symmetry algebra. Ultimately, the operator product algebra takes the following form:
@@
$\mathfrak{\mathcal{\mathfrak{J}}}^{A}(\tau_{1})\mathfrak{\mathcal{\mathfrak{J}}}^{B}(\tau_{2})$ $\sim$ $\frac{\frac{k}{2}\eta^{AB}}{\tau_{12}^{2}}+\frac{\mathfrak{\mathcal{\mathfrak{f}}}^{AB}_{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ C}\mathfrak{\mathcal{\mathfrak{J}}}^{C}}{\tau_{12}}.$ (67)
Here, $\eta^{AB}$ denotes the invariant trace matrix, while $\mathfrak{f}^{AB}_{\leavevmode\nobreak\ \leavevmode\nobreak\ C}$ represent the structure constants of the $\mathfrak{osp}(1|2)$ algebra, with indices $(A,B=0,\pm 1,\pm\tfrac{1}{2})$. These constants satisfy the relation $\mathfrak{f}^{ij}_{\leavevmode\nobreak\ \leavevmode\nobreak\ i+j}=(i-j)$. Under the most general (least restrictive) boundary conditions, the asymptotic symmetry algebra of $\mathfrak{osp}(1|2)$ dilaton supergravity is characterized by a single copy of the affine $\mathfrak{osp}(1|2)_{k}$ algebra.
@@ 4.2 $\mathcal{N}=1$ superconformal boundary conditions
These operator product expansions define the classical $\mathcal{N}=1$ superconformal algebra, which extends the Virasoro algebra through the inclusion of the spin-$\tfrac{3}{2}$ current $\mathcal{G}$. In particular, the operator product (79) captures the interaction between the spin–2 and spin–$\tfrac{3}{2}$ modes.
@@
From a physical perspective, the $\mathcal{N}=1$ superconformal algebra governs the dynamics of spin–$2$ and spin–$\tfrac{3}{2}$ boundary currents. The generator $\mathcal{L}$ corresponds to the energy–momentum tensor, while $\mathcal{G}$ represents a conserved higher–spin current of conformal weight–$\tfrac{3}{2}$. In the holographic setting, nontrivial $\mathcal{N}=1$ superconformal charges reflect the presence of spin–$\tfrac{3}{2}$ hair at the $\mathrm{AdS}_{2}$ boundary, leading to higher–derivative deformations in the dual boundary theory. The incorporation of the dilaton fields $\mathcal{X}$ and $\mathcal{Y}$ modifies the standard $\mathcal{N}=1$ algebra by introducing symmetry-breaking terms, thereby revealing how higher–spin symmetries become partially broken due to boundary conditions. This deformation parallels the symmetry-breaking mechanism observed in the Schwarzian limit of $sl(2,\mathbb{R})$ JT gravity, now generalized to spin-$\tfrac{3}{2}$ dynamics.
It follows that $\mathcal{X}$ and $\mathcal{Y}$ behave as boundary vectors, while $\mathcal{L}$ and $\mathcal{G}$ function as spin–$2$ and spin–$\tfrac{3}{2}$ densities, respectively. In this work, we have reformulated the boundary conditions in such a way that eliminates any explicit dependence on the gauge parameter $\lambda$. Consequently, the residual symmetries associated with $a_{t}$ and $x$ correspond to circle diffeomorphisms generated by the vector fields $\epsilon$ and $\eta$. For the conformal boundary conditions in $\mathfrak{osp}(1|2)$ dilaton supergravity, the resulting asymptotic symmetry algebra is identified as a single copy of the $\mathcal{N}=1$ $\mathfrak{osp}(1|2)$ superconformal algebra with central charge $c=3k$.

### v2

@@ 1 Introduction
The holographic principle Maldacena:1997re asserts that a gravitational theory defined in $d$ dimensions can be reformulated through an equivalent boundary theory living in (d1)(d{-}1) dimensions. This paradigm inspired investigations into two–dimensional $\mathrm{AdS}_{2}$ holography in the late 1990s, primarily driven by the analysis of asymptotic symmetries in Jackiw–Teitelboim (JT) gravity Teitelboim:1983ux ; Jackiw:1984je and dilaton models arising from dimensional reduction Strominger:1998yg ; Cadoni:1998sg ; Hotta:1998iq ; Navarro-Salas:1999zer . These studies have recently experienced a resurgence, motivated by the conjectured duality between JT gravity and the Sachdev–Ye–Kitaev (SYK) model Maldacena:2016hyu ; Jensen:2016pah ; Sachdev:1992fk ; Kitaev , which exemplifies a solvable system with quantum chaotic behavior. Within this correspondence, the Schwarzian boundary action emerges as a low-energy effective description Mertens:2018fds .
Conformal field theory has emerged as a foundational element in contemporary theoretical physics, with far–reaching applications from string theory to the study of complex many–body systems. Its significance lies in the stringent symmetry principles it enforces, which serve to tightly constrain the landscape of admissible physical models. In two dimensions, conformal invariance proves exceptionally potent, manifesting in the form of the infinite–dimensional Virasoro algebra Virasoro:1969zu . Over time, extensive research has been directed toward extending this symmetry to more generalized algebraic settings, such as the Kac–Moody algebras Kac_1968 ; Moody:1968zz and WW–algebras Zamolodchikov:1985wn . Simultaneously, substantial effort has been invested in deepening the theoretical understanding of the Virasoro structure itself DiFrancesco:1997nk ; Blumenhagen:2009zz ; Prochazka:2014gqa ; Ozer:2015dha . These algebraic symmetries also underpin the formulation of two–dimensional dilaton gravity, wherein the appearance of the Virasoro algebra at the boundary offers a natural bridge between conformal techniques and gravitational dynamics.
@@
The intricate internal structure of JT gravity is fundamentally grounded in its boundary symmetry content and the formal mathematical scaffolding that supports it. The theory maintains a profound connection to conformal symmetry through its interplay with the Virasoro algebra and Schwarzian derivatives Iliesiu ; Mertens ; OvisenkoSchwarzian , and it has been systematically broadened by the implementation of relaxed boundary conditions and higher–spin extensions Gross:2016kjj ; Witten:2016iux ; Yoon . The boundary behavior can also be effectively captured using coadjoint orbit methods, which serve as powerful tools for probing symmetry–breaking phenomena and their physical ramifications Davison:2016ngz ; NarayanYoon ; Fu:2016vas ; Grumiller:2017qao ; Alkalaev:2013fsa ; Grumiller:2013swa ; Grumiller ; Kine ; Fukuyama . While the Schwarzian formalism remains a crucial link to the SYK dual, the scope of asymptotic symmetry analysis extends beyond it; in particular, the BF formulation of JT gravity offers a structurally independent yet equally robust avenue for such inquiries Grumiller:2002nm ; Banks:1990mk . Within this formalism, the dilaton field functions as a Lagrange multiplier that enforces curvature constraints, and the associated asymptotic symmetry algebra can be extended, for instance, to a Virasoro–Kac–Moody structure, via the inclusion of internal symmetry sectors Ikeda:1993fh ; Jackiw:1982hg . This flexibility underscores the utility of JT gravity as a versatile platform for studying gravitational dynamics in lower dimensions, irrespective of adherence to the Schwarzian perspective. This methodological structure aligns with the strategy developed in our previous work on bosonic $\mathfrak{sl}(3,\mathbb{R})$ JT gravity Ozer:2025bpb , and here we implement its supersymmetric analogue using the $\mathfrak{osp}(1|2)$ superalgebra. While prior analyses of JT supergravity, particularly those centered around the Schwarzian action Almheiri:2014cka ; Maldacena:2016upp ; Cadoni:1999ja ; Engelsoy:2016xyb ; Cvetic:2016eiv ; Sachdev:2019bjn ; Sachdev:2010um , focus on the boundary effective theory and its relation to the SYK model, our approach emphasizes the bulk gauge-theoretic formulation and its implications for asymptotic symmetry breaking. In contrast to the boundary-centered derivations of the ${\cal N}=1$ super–Schwarzian theory, we construct the full $\mathfrak{osp}(1|2)$ asymptotic algebra via BF theory and coadjoint orbit methods, thereby uncovering dynamical constraints absent in Schwarzian-only treatments. This strategy is conceptually aligned with our earlier $\mathfrak{sl}(3,\mathbb{R})$ construction Ozer:2025bpb , and extends the findings of Grumiller:2002nm ; Banks:1990mk ; Ikeda:1993fh ; Jackiw:1982hg into the supersymmetric domain. As such, the present work provides a structurally distinct extension of JT supergravity beyond existing formulations.
@@
In this regard, the present construction is best interpreted as a supersymmetric extension of our previous bulk–based asymptotic symmetry analysis, rather than as a novel dynamical theory beyond the Schwarzian paradigm.
Our construction differs conceptually and technically from conventional JT gravity formulations based on $\mathfrak{sl}(2,\mathbb{R})$ Momeni:2024ygv ; Momeni:2020tyt ; Momeni:2020zkx , particularly those that emphasize boundary actions, dilaton constraints, or effective Schwarzian descriptions. While those models often encode symmetry breaking through boundary terms and thermodynamic consistency conditions, our approach derives asymptotic symmetries directly from the bulk via the structure of the gauge field and its residual supersymmetries. This method bypasses the need for an explicit boundary action, allowing a direct algebraic characterization of the asymptotic symmetry algebra. In that sense, our work aligns more closely with generalized bulk–based symmetry analyses, such as our earlier $\mathfrak{sl}(3,\mathbb{R})$ study Ozer:2025bpb , and represents a supersymmetric extension of that framework. We do not aim to construct a complete dual quantum mechanical theory; rather, we focus on the intrinsic gauge–theoretic origin of boundary symmetries in the supersymmetric BF setting.
Unlike earlier JT models based on $\mathfrak{sl}(2,\mathbb{R})$, where symmetry breaking typically appears through boundary dilaton constraints or horizon regularity conditions, our formulation realizes symmetry reduction algebraically via residual bulk supersymmetry, without imposing explicit dynamical constraints or thermodynamic inputs. On the physical level, symmetry breaking happens dynamically due to the time-dependent dilaton supermultiplet in the bulk, not because of boundary thermodynamics.
@@ 2.1 Conformal structure of the quantum ${\cal N}=1$ superconformal algebra
[…]
In this section, we construct an extended version of the quantum ${\cal N}=1$ superconformal algebra by supplementing the conformal fields of spin $\Delta=2$ and $\Delta=3/2$, introduced in the previous subsection, with two additional components. These new fields, labeled $X$ and $Y$, are interpreted as dilaton partners and carry conformal weights $1-\Delta=-1$ and 1/2-1/2, respectively. Their standard mode expansions are given by $X(z)=\sum_{n}X_{n}z^{-n+1}$ and $F(z)=\sum_{n}F_{n}z^{-n+1/2}$. Using Thielemans’ Mathematica package Thielemans:1991uw , one obtains the non–vanishing operator product expansions that define the extended ${\cal N}=1$ superconformal algebra in this formulation.
@@ 3.2 Connection to $sl(2,\mathbb{R})$ BF theory
[…]
@@
[…]
@@
$\operatorname{d}\!s^{2}=\operatorname{d}\!\rho^{2}+2\,\mathcal{L}^{0}\operatorname{d}\!\rho\operatorname{d}\!\varphi+\Big(\left(e^{\rho}\mathcal{L}^{+}-e^{-\rho}\mathcal{L}^{-}\right)^{2}+\left(\mathcal{L}^{0}\right)^{2}\Big)\,\operatorname{d}\!\varphi^{2}\penalty 10000\ ,$ (18)
@@
$X=e^{\rho}\,\mathcal{X}^{+}+e^{-\rho}\,\mathcal{X}^{-}\penalty 10000\ .$ (19)
@@ 3.3 Affine boundary conditions
[…]
@@
$\mathfrak{\mathcal{\mathfrak{J}}}^{A}(\tau_{1})\mathfrak{\mathcal{\mathfrak{J}}}^{B}(\tau_{2})$ $\sim$ $\frac{\frac{k}{2}\eta^{AB}}{\tau_{12}^{2}}+\frac{\mathfrak{\mathcal{\mathfrak{f}}}^{AB}_{\penalty 10000\ \penalty 10000\ \penalty 10000\ C}\mathfrak{\mathcal{\mathfrak{J}}}^{C}}{\tau_{12}}.$ (33)
Here, $\eta^{AB}$ denotes the trace matrix, and $\mathfrak{f}^{AB}_{\penalty 10000\ \penalty 10000\ C}$ are the structure constants of the underlying algebra, with indices running over $(A,B=0,\pm 1)$. In particular, one has $\mathfrak{f}^{ij}_{\penalty 10000\ \penalty 10000\ i+j}=(i-j)$. Accordingly, under the most relaxed set of boundary conditions for $sl(2,\mathbb{R})$ dilaton gravity, the resulting asymptotic symmetry algebra is identified with a single copy of the affine $\mathfrak{sl}(2,\mathbb{R})_{k}$ algebra.
@@ 3.4 Conformal boundary conditions
[…]
@@ 4.1 $\mathcal{N}=1$ affine boundary conditions
$\delta_{\lambda}\mathcal{G}^{\mp\frac{1}{2}}$ $=\pm k\partial_{t}\eta^{\mp\frac{1}{2}}+2\eta^{\pm\frac{1}{2}}\mathcal{L}^{\mp 1}+2\eta^{\mp\frac{1}{2}}\mathcal{L}^{0}-\mathcal{G}^{\pm\frac{1}{2}}\epsilon^{\mp 1}\mp\frac{1}{2}\mathcal{G}^{\mp\frac{1}{2}}\epsilon^{0},$ (53)
@@
$\delta_{\lambda}\mathcal{Y}^{\mp\frac{1}{2}}$ $=\pm\frac{1}{2}\eta^{\mp\frac{1}{2}}\mathcal{X}^{0}\mp\eta^{\pm\frac{1}{2}}\mathcal{X}^{\mp 1}\pm\mathcal{Y}^{\pm\frac{1}{2}}\epsilon^{\mp 1}\mp\frac{1}{2}\mathcal{Y}^{\mp\frac{1}{2}}\epsilon^{0}.$ (56)
@@
[…]
@@
$\mathfrak{\mathcal{\mathfrak{J}}}^{A}(\tau_{1})\mathfrak{\mathcal{\mathfrak{J}}}^{B}(\tau_{2})$ $\sim$ $\frac{\frac{k}{2}\eta^{AB}}{\tau_{12}^{2}}+\frac{\mathfrak{\mathcal{\mathfrak{f}}}^{AB}_{\penalty 10000\ \penalty 10000\ \penalty 10000\ C}\mathfrak{\mathcal{\mathfrak{J}}}^{C}}{\tau_{12}}.$ (67)
Here, $\eta^{AB}$ denotes the invariant trace matrix, while $\mathfrak{f}^{AB}_{\penalty 10000\ \penalty 10000\ C}$ represent the structure constants of the $\mathfrak{osp}(1|2)$ algebra, with indices $(A,B=0,\pm 1,\pm\tfrac{1}{2})$. These constants satisfy the relation $\mathfrak{f}^{ij}_{\penalty 10000\ \penalty 10000\ i+j}=(i-j)$. Under the most general (least restrictive) boundary conditions, the asymptotic symmetry algebra of $\mathfrak{osp}(1|2)$ dilaton supergravity is characterized by a single copy of the affine $\mathfrak{osp}(1|2)_{k}$ algebra.
@@ 4.2 $\mathcal{N}=1$ superconformal boundary conditions
[…]
@@
To illustrate the mechanism of symmetry breaking induced by the dilaton supermultiplet, we present a simplified toy model based on a static, partially truncated configuration. Consider the following profile for the $\mathfrak{osp}(1|2)$–valued dilaton field:
$\mathcal{X}^{0}=\text{const},\quad\mathcal{X}^{\pm 1}=0,\quad\mathcal{Y}^{\pm\frac{1}{2}}=0.$
This choice retains only the $\mathcal{X}^{0}$ component while suppressing all fermionic modes $\mathcal{Y}^{\pm\frac{1}{2}}$ and the bosonic charges $\mathcal{X}^{\pm 1}$. Under the gauge transformations generated by the full affine algebra (Eqs. (51)–(56)), one finds that such a configuration leads to a restricted set of residual symmetries. In particular, the variations $\delta_{\lambda}\mathcal{Y}^{\pm\frac{1}{2}}$ and $\delta_{\lambda}\mathcal{X}^{\pm 1}$ vanish identically, while $\delta_{\lambda}\mathcal{X}^{0}$ survives only for $\epsilon^{0}$ and a subset of $\eta$–parameters. As a consequence, the boundary charges associated with broken modes cease to transform non–trivially and are effectively projected out of the asymptotic symmetry algebra. This demonstrates, at a minimal level, how specific dilaton backgrounds can dynamically suppress certain symmetry generators, resulting in a reduced realization of the full $\mathfrak{osp}(1|2)_{k}$ structure. This kind of toy model construction, where a specific field configuration is used to analyze residual symmetry algebras, is also employed in the context of AdS2 boundary conditions Grumiller:2017qao .
@@
It is important to emphasize that the conformal boundary conditions considered in this section are not merely a technical simplification, but rather encode the infrared dynamics of the boundary theory. While the affine boundary data support a richer set of boundary modes—including commuting currents and higher–spin extensions—the conformal reduction effectively isolates the pseudo–Goldstone mode associated with the breaking of reparametrization symmetry. This reduction aligns with the low–energy behavior of dual SYK–like systems and captures the dominant physical degree of freedom via the Schwarzian effective action. Therefore, the transition from affine to conformal boundary conditions should be viewed as a flow from a general UV–complete description to an IR–dominated effective theory.
4.3 Scope and physical context
The present work is conceived as the $\mathcal{N}=1$ supersymmetric extension of our earlier analysis of $\mathfrak{sl}(3,\mathbb{R})$ JT–gravity Ozer:2025bpb . In that framework, asymptotic symmetry structures were studied from a bulk perspective, without recourse to Schwarzian–type boundary actions. The method focused on identifying residual gauge transformations and symmetry breaking patterns arising directly from the bulk gauge fields, providing a non–boundary formulation of the ASA.
In this supersymmetric extension, the same philosophy is maintained: we do not attempt to introduce a new effective boundary action, nor do we claim to formulate an alternative physical regime. Rather, we aim to generalize the bulk gauge–theoretic approach to the supersymmetric case, governed by the $\mathfrak{osp}(1|2)$ superalgebra.
The phrase “beyond the Schwarzian regime” is therefore not intended to suggest a new physical theory, but simply refers to a methodological shift away from boundary action–based analyses. This approach is rooted in the structure of gauge symmetries and their asymptotic realization, consistent with the strategy we established in the non–supersymmetric ${\tt S}{\tt L}(3,\mathbb{R})$ setting.
In this sense, while our construction remains algebraic in scope, the extended bulk symmetry breaking mechanism within the $\mathfrak{osp}(1|2)$ framework may serve as a foundation for future developments involving boundary dynamics, supersymmetric observables, or holographic models that transcend the traditional Schwarzian description. We believe that this perspective may complement ongoing efforts to understand JT supergravity beyond its standard boundary effective action paradigm.
From a physical standpoint, the dilaton-induced symmetry breaking mechanism analyzed in this work can be expected to influence observables in putative holographic dual theories. Specifically, the reduction of the full affine $\mathfrak{osp}(1|2)_{k}$ symmetry to its residual finite-dimensional subalgebra—together with the emergence of abelian commuting modes—modifies the spectrum of conserved charges at the boundary. This structure suggests a constrained class of boundary excitations, possibly leading to nontrivial selection rules or effective couplings in dual quantum mechanical models. While an explicit construction of the dual theory lies beyond the scope of this paper, we anticipate that the algebraic features derived here may provide structural constraints on supersymmetric SYK–like models or other one–dimensional superconformal systems. Further investigation of these physical implications remains an important direction for future work.
We conclude this section by briefly commenting on the quantum aspects alluded to above. While the central extension of the superalgebra emerges naturally from the classical coadjoint orbit analysis, potential operator ordering ambiguities can affect the precise realization of the symmetry generators at the quantum level. Addressing such ambiguities may require a more systematic treatment—such as path integral quantization or BRST analysis—which lies beyond the scope of this work. Nonetheless, our current results offer a robust classical foundation for such future quantum investigations, particularly in supersymmetric generalizations of the Schwarzian theory.
As a future direction, it would be interesting to explore quantum extensions of the symmetry breaking mechanisms presented here, especially in the context of super–Schwarzian dynamics or SYK–like boundary theories. The toy model introduced in Section 4.2 may also provide a useful testing ground for probing correlation functions and operator spectra in dual descriptions. We hope that our results provide a useful platform for further developments in two–dimensional supergravity and its holographic applications.
Quantum Extensions and Ordering Effects.
Beyond the classical analysis, the incorporation of quantum aspects introduces subtleties that merit closer attention. One such feature is the appearance of central extensions, which in our setup arise naturally from the coadjoint orbit construction. In particular, the Kirillov–Kostant symplectic structure on the orbit allows for a well-defined symplectic form, from which the classical central term in the $\mathcal{N}=1$ superconformal algebra can be read off Grumiller:2017qao ; Ikeda:1993fh . However, when promoting this classical structure to the quantum level, further issues emerge.
Most notably, operator ordering ambiguities may affect the precise realization of the symmetry generators and their algebra. These ambiguities stem from the non-commutative nature of quantum operators, especially in systems with fermionic generators and nontrivial constraints, such as in the $\mathfrak{osp}(1|2)$ case. While the classical Poisson brackets yield a consistent algebra, the quantum commutators require careful normal ordering or regularization prescriptions to preserve associativity and supersymmetry.
To properly resolve these quantum-level challenges, several quantization approaches can be considered:
  • Coadjoint orbit quantization: constructs the Hilbert space by quantizing the orbit equipped with a symplectic form Grumiller:2017qao ; Ikeda:1993fh . This technique is geometrically transparent but faces challenges in the presence of superalgebraic structure and boundary conditions.
  • BRST quantization: systematically incorporates gauge symmetries and constraints into the quantum framework Ikeda:1993fh ; Grumiller:2013swa . It is particularly suitable for supergravity models and BF-type gauge theories but often trades off geometric intuition for algebraic rigor.
  • Path integral quantization: provides a functional integral formulation of the theory, enabling the study of partition functions and quantum correlation functions Engelsoy:2016xyb ; Cardenas:2018krd . For $\mathfrak{osp}(1|2)$ supergravity, this approach captures the boundary dynamics and fluctuations of super-Schwarzian modes.
While a detailed quantum treatment lies beyond the present scope, the formal framework established in this work offers a solid foundation for such investigations. In particular, the toy model and the algebraic symmetry breaking discussed in Section 4.2 provide concrete starting points for analyzing quantum deformations and the structure of dual one-dimensional superconformal systems.